# Congestion-Aware Path Re-routing Strategy for Dense Urban Airspace

### v1

@@ V Theoretical and Simulation Results
Figure 19 shows snapshots of UAS distributedly employing the proposed rule-based congestion mitigation strategy. When every UAS in the source-destination traffic stream follows the proposed strategy, Fig. 20-22 show that the emerging traffic pattern closely resembles a parallel air corridor network with the number of air corridors adapting to the arrival rate  [^1] The simulation videos have been uploaded in the following link https://youtu.be/Yk1vY6nynHg. The emergence of parallel air corridors is a consequence of the proposed strategy and UTM has no role in it nor there is any scheduling policy in place. It may be observed that in any zone no more than $M$ UAS are present, thus assuring increased UAS intrinsic safety. The above figures also demonstrate the variation in the number of active corridors, also referred to as traffic stream spread as a function of the arrival rate ($\lambda$) and the UTM design parameters ($M,\ \eta$). As the arrival rate increases or $M$ decreases, the congestion frequency and duration in the inner stream zones increases, hence more UAS shift onto outer streams, resulting in increased traffic spread. When $\eta<0.5$ we have denser traffic flow on the left of the nominal segment and vice versa.
When the arrival rate is small, in the lower levels the UAS inter-separation distances (depend only on arrival) are large. Hence the congestion duration is small (refer Fig. 23a). However in the higher levels, when a zone in the inner stream turns uncongested, the UAS in outer streams descend onto the inner stream (UAS marked 1 in snapshot 8 of Fig. 19). Thus, reducing UAS inter-separation distances, and resulting in longer congestion durations (refer Fig. 23b). In the $Z=(0,1)$ zone, when $\lambda=1$, a UAS is available to enter every timeslot the zone turns uncongested (refer Fig. 23c). Thus, the zone remains congested for $S-M$ timeslots. For example, when $S=11,\ M=2$ and $\lambda=1$, the zone is congested for $S-M=9$ timeslots and hence the congestion probability $\Theta_{0}=(S-M)/S=0.82$ (as validated in the $\Theta_{0}$ plot of Fig. 24, when $\lambda=1$). The rearrangement of UAS in the traffic stream happens such that beyond a certain level, we observe no UAS moving outward and no UAS descending inward.
@@
Figure 19: The UTM parameters are $M=3$ and $\eta=0.5$. The arrival rate $\lambda$ at source is $0.2$. In snapshot 1, three $L=5$ slot look-ahead positions are present in $Z_{\mathrm{UP}}$. As $M=3$, from the perspective of the UAS present in $Z$, the zone $Z_{\mathrm{UP}}$ is congested (Rule 1). In snapshots 2-3, when UAS 1, 2 arrived the $Z_{\mathrm{UP}}$ was congested, hence these UAS entered the queue (Rule 2). In snapshot 3, as UAS in the $\gamma\rightarrow\beta$ path (UAS 4) are excluded in the congestion definition, only two look-ahead positions are in $Z_{\mathrm{UP}}$. In snapshots 4-5, as $Z_{\mathrm{UP}}$ is uncongested, the UAS 2, 3 enter service following the LCFS discipline (Rule 3). As there is a queuing deadline of $L=5$ slots, UAS 1 is lost from the $\overline{Z}$ queuing system. In snapshot 6, the past trajectory of UAS 1,2,3 shows the path opted corresponding to $\psi>0$, $\psi<0$ and $\psi=0$, respectively (Rule 4). In snapshot 7-8, from the perspective of UAS 1, as $Z_{\mathrm{UD}}$ is uncongested and no UAS are present in the $\gamma_{Z}\rightarrow\beta_{Z}$ path, the descend condition is satisfied. Hence UAS 1 descends to Stream 0 (Rule 6). Whereas, in snapshots 9-10, though $Z_{\mathrm{UD}}$ is uncongested the descend condition is not satisfied because of UAS 6, hence UAS 5 does not descend.
@@
Figure 21: Variation of UAS traffic stream spread as a function of bias $\eta\in[0.1,0.3,0.5,0.7,0.9]$ in traffic distribution about Stream 00 segment when $M=3$ and $\lambda=1.0$.
@@
Figure 22: Variation of UAS traffic stream spread as a function of the minimum number of UAS to be present in a zone for congestion $M\in[2,3,4]$ when $\eta=0.5$ and $\lambda=0.8$. In any zone, at most $M$ number of UAS are present.
The same is validated in Fig. 24, the average number of UAS in the queue (that is UAS moving outward), and the queue overflow is nearly zero for Level $2$ and above zones. In the congestion probability plot, when $\lambda<0.2$, there is no congestion, so no UAS shifts onto outer streams and hence no UAS available to descend. The inter-separation distances depend only on arrival. Thus, the congestion duration remains unaffected as levels increase. When $\lambda\in(0.2,0.5)$, there is congestion but the duration is small. There are UAS that descend and increase the congestion duration in the higher levels. However, on average the increase is not significant as with the increase in the congestion duration, the probability that a UAS would descend decreases. When $\lambda>0.6$, the congestion durations are long enough that most of the UAS arriving in the lower levels shift onto outer streams. However, these UAS rarely descend (refer Fig. 25b). Hence again the congestion duration remains unaffected as the level increases. Thus as $\lambda\rightarrow 0$ and as $\lambda\rightarrow 1$ the congestion probabilities of all $(0,Y)$ zones converge.
Unlike arrival rate $\lambda$ which cannot be controlled, the UTM can take strategic decisions for parameters $M$ and $\eta$. Figure 26 shows Rule 7 being employed by UAS in the presence of a static obstacle, due to which the UAS circumnavigate the obstacle. Instead, the UTM could choose an $\eta$ such that the traffic spread (shown in Fig. 21) would not pass through the static obstacle. In general, the traffic flow in the urban airspace is weather-dependent. As per UAS safety standards, the allowable congestion in a region is constrained on wind conditions. As in [19], the weather is modeled as a first-order Markov chain with zero, moderate, and heavy wind as states. The UTM may associate each state with a $M$ value (shown in Fig. 27a). Corresponding to prevailing wind condition, the UTM communicates respective $M$ value with the UAS. The UAS accordingly adapt ensuring no more than $M$ UAS are present in any zone, as shown in Fig. 22. Another use case for changing $M$ would be at the destination. $M$ must be gradually incremented over the levels (as shown in Fig. 27b) so that when $M=S$ in the final level containing the destination node, all UAS would enter the destination.
@@
Figure 23: The congestion duration when $M=2$ and a) $\lambda$ is small, b) $\lambda$ is small but there are UAS in the outer stream that would descend, c) $\lambda=1$.
Figure 24: Figure presents theoretical and simulation comparison of congestion in $Z=(1,Y)$ zone, $Y=\mathbb{Z}_{[1,10]}$ for varying arrival rate $\lambda\in[0.05,1]$ and minimum number of UAS in zone for congestion $M=2$. The UTM parameters are $S=2L+1=11$ cells, and $\eta=0.5$.
@@
Figure 25: When $M=2$ and $\eta=0.5$, the figure shows a) the time delay introduced in the UAS path as a consequence of the UAS executing lateral transitions in the event of congestion, and b) the number of upstream and diagonal transitions executed by the UAS to reach a Level $10$ zone (that is a total of $11010\times S=$ 110 transitions).
@@
When executing a diagonal transition, to travel 1 unit distance in the upstream direction the UAS needs $\sqrt{2}$ times higher velocity than when executing an upstream transition. Thus, diagonal transitions are energy-intensive maneuvers. However, as shown in Fig. 25b, the number of diagonal transitions executed is low and mostly confined to the lower levels.
@@
In Fig. 26a, the UAS marked * in zone $1$ always perceive zone $1_{\mathrm{UP}}$ to be congested. These UAS would certainly enter zone $Z_{\mathrm{UP}}$. The respective look-ahead positions can be determined and hence these UAS are included in Definition 2. Thus, UAS in zone $Z$ are affected by UAS *.
@@
Figure 27: Figure presents use cases for controlling UTM parameter $M$, a) weather-dependent congestion regulation where wind conditions are modeled as a Markov chain and b) UAS scheduling at the destination.
Figure 28: Figure presents theoretical and simulation comparison of congestion for $L=5,\ S=2L+1=11$ and varying arrival rate $\lambda\in[0.05,1]$, in a) Stream 00 Level $1$ zone when the minimum number of UAS in the zone for congestion $M\in[2,4,6,8,10]$; and in b, c) Stream $-1$ Level $1$ zone when $M\in[2,3,4,5,6]$ and $\eta=0.5$ and $\eta=0$.
@@
For any given $L,\ M$ and $\eta$, the UTM could estimate the expected traffic spread, and the expected number of UAS in the zone beforehand using the queueing models discussed in section IV. Based on these estimates the UTM could decide on the look-ahead window $L$ to define the zone and then decide on the parameters $M$ and $\eta$ for the appropriate use case. Fig 28-29 provide validation for theoretical estimates obtained from the queuing models.
@@
When $\lambda=1$, we have an arrival every timeslot in the $Z=(0,1)$ zone. As mentioned before, the zone would be congested for $S-M$ timeslots, and hence congestion probability $\Theta_{0}$ is $(S-M)/S$. The same can be verified in Fig. 28a for different values of $M$. The average number of UAS in service never exceeds $M$. The UAS queuing would be in either of the two $\beta_{0}$ queues each of capacity $L$, with probability $\eta$ and ($1-\eta$). Hence, the average number of UAS in the queue never exceeds $L$. In Fig. 28-29 for zone $Z=(\pm X,1),\ X\neq 0$, there maybe UAS queuing in both $\gamma$ and $\beta$ queues. Hence, the average number of UAS queueing never exceeds $12L-$ 1. Only a fraction of the UAS arrivals to zone $Z=(0,1)$ are lost from the $\beta_{0}$ queue to become non-nodal arrivals for $Z=(\pm 1,1)$ zones. For a given arrival rate $\lambda$, the zone $(0,1)$ has higher congestion probability $\Theta_{0}$ than zone (±$(\pm 1,1)$ as less number of UAS are available in (±$(\pm 1,1)$ zones to enter service. For the same reason, zones $(X,1)$ and $(-X,1)$ have greater $\Theta_{0}$ than outer neighboring $(X+1,1)$ and $(-X-1,1)$ zones, respectively. When $\eta=0.5$, we have equal $\Theta_{0}$ in (±$(\pm X,1)$ zones. When $\eta=0$ all UAS lost from $Z=(0,1)$ become non-nodal arrivals to $Z=(-1,1)$ else only $(1-\eta)$ fraction of UAS lost become non-nodal arrivals. For any given $\lambda$, as more number of UAS are available to enter service when $\eta=0$, hence in $Z=(-1,1)$ we have more number of UAS in service, resulting in increased congestion, and more number of UAS in the queue when compared to $\eta>0$.
@@
V-A Presence of exogenous traffic stream
@@
Following Assumptions 1-3, an exogenous traffic stream perpendicularly intersecting the nominal segment at Level 2 is considered (refer Fig. 30a). The exogenous UAS arrival rate is $\lambda_{e}=0.2$. The expected number of exogenous UAS in the Level $2$ zones is $S\lambda_{e}=2.2$. $M$ must be greater than $2.2$ for UAS to cross the exogenous traffic stream, hence $M$ is chosen to be $4$. The exogenous UAS transitions with $1$ cell per time slot nominal velocity. The exogenous UAS do not follow the proposed mitigation strategy and are hence unaffected by congestion in the zone. However, the exogenous UAS contributes to congestion. When $M$ or more than $M$ look-ahead positions (exogenous UAS included) are present in a zone, then the UAS laterally transitions to avoid such zone (as shown in snapshot 1 of Fig. 30a). Unlike in the presence of a static obstacle, where it is certain that a UAS would always laterally transition every timeslot when in $\gamma\rightarrow\beta$ path, in the presence of an exogenous stream when the UAS would laterally transition is uncertain (congestion-dependent) and thus are not included in Definition $2$. This results in the UAS transitioning laterally and those transitioning upstream experiencing conflicts (as shown in snapshot 1-2 of Fig. 30a). A conflict among UAS belonging to the same stream is referred to as internal conflict. A conflict between a UAS and an exogenous UAS is referred to as exogenous conflict. Though conflicts occur, the congestion mitigation strategy guarantees no more than $M$ conflicts in any zone, that is UAS intrinsic safety is improved. Further, the mitigation strategy helps reduce the total number of conflicts in the stream (as shown in Fig. 30b). These advantages come at the expense of additional time delay in the UAS path and energy cost proportional to the number of diagonal transitions (refer Fig. 30b. Here, the total number of upstream and diagonal transitions executed to reach Level $5$ zone is $555\times S=$ 55). Note that the additional cost is insignificant when computed over the entire flight path. Depending on the inter-separation distance distribution between the exogenous UAS, the traffic spread may increase or decrease as shown in snapshots 3-4 of Fig. 30a. However, the queueing models can be used to estimate the expected traffic spread as validated in Fig. 30c.
@@
Figure 30: Congestion mitigation being employed by UAS in the presence of an exogenous traffic stream when $M=4$ and $\eta=0.3$. a) As a consequence, in snapshot 1, UAS 1 conflicts with UAS 2 belonging to the same stream at timeslot 47, and at timeslot 49 UAS 3 would conflict with an exogenous UAS 4. Rule 5 is employed by UAS to resolve the conflict as shown in snapshot 2, b) Comparison of the number of conflicts with ($M=4$) and without ($M=S$) the proposed congestion mitigation strategy and trade-off concerning time-delay and energy cost associated with the maneuver. c) Validation of expected traffic stream spread estimation in the presence of an exogenous stream ($\lambda_{e}=0.2$).
VI Conclusion

### v2

@@ 5 Simulation Results and Analysis
Every UAS deployed at the source follows the proposed congestion mitigation strategy. The simulation is run until around $1200$ UAS are deployed at the source. The congestion probability and expected number of UAS present in the zone are numerically computed by solving Eqn. 11 using MATLAB symbolic math toolbox. The theoretical results are validated with the average values of the $1200$ UAS simulation data.
@@
Figure 16 shows snapshots of UAS employing the proposed rule-based congestion mitigation strategy in a distributed manner. When every UAS in the source-destination traffic stream follows the proposed strategy, Fig. 17-19 show that the emerging traffic pattern closely resembles a parallel air corridor network with the number of air corridors adapting to the arrival rate  [^1] The simulation videos have been uploaded in the following link https://youtu.be/Yk1vY6nynHg. The emergence of parallel air corridors is a consequence of the proposed strategy, and UTM has no role in it, nor is there any scheduling policy in place. It may be observed that in any zone no more than $M$ UAS are present, thus assuring increased UAS intrinsic safety. The above figures also demonstrate the variation in the number of active corridors (also referred to as traffic stream spread) as a function of the arrival rate ($\lambda$) and the UTM design parameters ($M,\ \eta$).
@@
Figure 16: The arrival rate $\lambda$ at source is $0.2$. The UTM parameters are $M=3$ and $\eta=0.5$. In snapshot 1, three $L=5$ slot look-ahead positions (that of UAS a,b,c) are present in $Z_{\mathrm{UP}}$. As $M=3$, from the perspective of the UAS present at the source node, the zone $Z_{\mathrm{UP}}$ is congested (Rule 1). In snapshot 2, when UAS 1 arrived, the $Z_{\mathrm{UP}}$ was congested; hence, UAS 1 entered the queue (Rule 2). In snapshot 3, the $L$-slot look-ahead position of UAS in the $\beta\rightarrow\gamma$ path (that of UAS 4) cannot be determined; only two look-ahead positions are in $Z_{\mathrm{UP}}$. In snapshots 3-5, as $Z_{\mathrm{UP}}$ is uncongested, the UAS 2, 3 enter service following the LCFS discipline (Rule 3). As there is a queuing deadline of $L=5$ slots, UAS 1 is lost from the $\overline{Z}$ queuing system. In snapshot 6, the past trajectory of UAS 1,2,3 shows the path opted corresponding to $\psi>0$, $\psi<0$ and $\psi=0$, respectively (Rule 4). In snapshot 7-8, from the perspective of UAS 1, as $Z_{\mathrm{ID}}$ is uncongested and no UAS are present in the $\beta_{Z}\rightarrow\gamma_{Z}$ path, the descend condition is satisfied. Hence, UAS 1 descends to $Z_{\mathrm{ID}}$ (Rule 5). Whereas, in snapshots 9-10, though $Z_{\mathrm{ID}}$ is uncongested, the descend condition is not satisfied because of UAS 6. Hence, UAS 5 does not descend.
@@
Figure 18: Variation of UAS traffic stream spread as a function of bias $\eta\in[0.1,0.3,0.5,0.7,0.9]$ in traffic distribution about the nominal segment when $M=3$ and $\lambda=1.0$. At most $3$ UAS are present in any zone.
@@
Figure 19: Variation of UAS traffic stream spread as a function of the minimum number of UAS to be present in a zone for congestion $M\in[2,3,4]$ when $\eta=0.5$ and $\lambda=0.8$. At most $M$ number of UAS are present in any zone. Snapshots 2 and 4 show the transient behavior of UAS moving outward and descending inward, respectively, adapting to the changing congestion parameter $M$.
As the arrival rate $\lambda$ increases or $M$ decreases, the congestion frequency and congestion duration in the inward stream zones increases. Hence, more UAS shift onto outward streams, resulting in increased traffic spread. When $\eta<0.5$, we have denser traffic flow on the left of the nominal segment and vice versa. In the lower levels, the UAS inter-separation distance depends only on the arrival rate. When the arrival rate is small, the UAS inter-separation distances are large. Hence, the congestion duration is small (refer Fig. 20a). However, in the higher levels, when a zone in the inward stream becomes uncongested, the UAS in outward streams descend onto the inward stream (UAS 1 in snapshot 8 of Fig. 16). Thus reducing UAS inter-separation distances and resulting in longer congestion durations (refer Fig. 20b). In the $Z=(0,1)$ zone, when $\lambda=1$, every timeslot the zone is uncongested, a UAS is available to enter the zone (refer Fig. 20c). Thus, the zone remains congested for $S-M$ timeslots. For example, when $S=11,\ M=2$ and $\lambda=1$, the zone is congested for $S-M=9$ timeslots and hence the congestion probability $\Theta_{0}=(S-M)/S=0.82$ (as validated in the $\Theta_{0}$ plot of Fig. 21, when $\lambda=1$). The rearrangement of UAS in the traffic stream happens such that beyond a certain level $Y$, we observe no UAS moving outward and no UAS descending inward.
@@
Figure 20: The congestion duration when $M=2$ and a) $\lambda$ is small, b) $\lambda$ is small, but there are UAS in the outward stream zone that would descend, c) $\lambda=1$, a UAS is available to enter every timeslot the zone is uncongested.
The same is validated in Fig. 21. The average number of UAS in the queue (that is, UAS moving outward) and the queue overflow is nearly zero for level $2$ and above zones. In the congestion probability plot, when $\lambda<0.2$, there is no congestion, so no UAS shifts onto outward streams, and hence, no UAS is available to descend. The inter-separation distances depend only on arrival. Thus, the congestion duration remains unaffected as levels increase. When $\lambda\in(0.2,0.5)$, there is congestion, but the duration is small. There are UAS that descend and increase the congestion duration in the higher levels. However, on average, the increase is not significant as with the increase in the congestion duration, the probability that a UAS would descend decreases. When $\lambda>0.6$, the congestion durations are long enough that most of the UAS arriving in the lower levels shift onto outward streams. However, these UAS rarely descend (refer Fig. 21c). Hence, again, the congestion duration remains unaffected as the level increases. Thus, as $\lambda\rightarrow 0$ and as $\lambda\rightarrow 1$, the congestion probabilities of all $(0,Y)$ zones converge.
@@
Figure 21: a) Theoretical and simulation comparison of congestion in $Z=(0,Y)$, $Y\in\{1,...,10\}$ for varying arrival rate $\lambda\in[0.05,1]$ and minimum number of UAS in zone for congestion $M=2$. Here, $S=2L+1=11$ cells, and $\eta=0.5$, b) The time delay introduced in the UAS path as a consequence of the UAS executing lateral transitions in the event of congestion and the number of upstream and, c) diagonal transitions executed by the UAS to reach a Level $10$ zone (that is a total of $11010\times S=$ 110 transitions).
@@
When executing a diagonal transition, to travel 1 unit distance in the upstream direction, the UAS needs $\sqrt{2}$ times higher velocity than when executing an upstream transition. Thus, diagonal transitions are energy-intensive maneuvers. However, the number of diagonal transitions executed is low (refer Fig. 21c) and mostly confined to the lower levels.
Unlike arrival rate $\lambda$, the UTM can take strategic decisions for parameters $M$ and $\eta$. Figure 22 shows Rule 7 being employed by UAS in the presence of a static obstacle, due to which the UAS circumnavigates the obstacle. Instead, the UTM could choose an $\eta$ such that the traffic spread would not pass through the static obstacle (as in Fig. 18). In Fig. 22, the UAS in zone $1$ always perceives zone $1_{\mathrm{UP}}$ to be congested. These UAS would certainly enter zone marked $2_{\mathrm{UP}}$. The respective $L$-slot look-ahead positions can be determined (indicated with blue dots in zone $2_{\mathrm{UP}}$) and contribute to congestion in Definition 2. Thus, the UAS in zone $2$ are affected by UAS laterally transitioning in the $\beta\rightarrow\gamma$ path of zone $1$. The traffic flow in the urban airspace is weather-dependent. As per UAS safety standards, the allowable congestion in a region is constrained by wind conditions. As in Zhou et al. (2020), the weather is modeled as a first-order Markov chain with zero, moderate, and heavy wind as states. The UTM may associate each state with a $M$ value (shown in Fig. 23a). Corresponding to prevailing wind conditions, the UTM communicates respective $M$ values with the UAS. The UAS accordingly adapts, ensuring no more than $M$ UAS are present in any zone, as shown in Fig. 19. Another use case for changing $M$ would be at the destination. $M$ must be gradually incremented over the levels (as shown in Fig. 23b) so that when $M=S$ in the final level containing the destination node, all UAS would enter the destination.
@@
Figure 23: Use cases for controlling UTM parameter $M$, a) weather-dependent congestion regulation where wind conditions are modeled as a Markov chain and b) UAS scheduling at the destination.
@@
For any given $L,\ M$ and $\eta$, the UTM could estimate the expected traffic spread and the expected number of UAS in the zone beforehand using the queueing models discussed in section 4. Based on these estimates, the UTM could decide on the look-ahead window $L$ to define the zone and then decide on the parameters $M$ and $\eta$ for the appropriate use case. Figures 24-25 provide validation for theoretical estimates obtained from the queuing models.
@@
When $\lambda=1$, there is an arrival entering the zone $Z=(0,1)$ in every timeslot. As mentioned before, the zone would be congested for $S-M$ timeslots, and hence congestion probability $\Theta_{0}$ is $(S-M)/S$. The same can be verified in Fig. 24a for $\lambda=1$ at different values of $M$. The average number of UAS in service never exceeds $M$. The UAS queuing would be in either of the two $\gamma_{0}$ queues, each of capacity $L$, with probability $\eta$ and ($1-\eta$). Hence, the average number of UAS in the queue never exceeds $L$. For zone $Z=(\pm X,1)$, where $X>0$, the UAS would be queuing in both $\gamma$ and $\beta$ queues. The average number of UAS queueing never exceeds $12L-$ 1. Only a fraction of the UAS arrivals to zone $Z=(0,1)$ are lost from the $\gamma_{0}$ queue to become non-nodal arrivals for $Z=(\pm 1,1)$ zones. For a given arrival rate $\lambda$, the zone $Z=(\pm 1,1)$ has less number of UAS available to enter service than zone $Z=(0,1)$. Hence, zone $Z=(\pm 1,1)$ has lower congestion probability $\Theta_{0}$ than zone $Z=(0,1)$. For the same reason, zones $(X+1,1)$ and $(-X-1,1)$ zones have lower $\Theta_{0}$ than inward neighboring zones $(X,1)$ and $(-X,1)$, respectively. When $\eta=0.5$, we have equal $\Theta_{0}$ in (±$(\pm X,1)$ zones. When $\eta=0$ all UAS lost from $Z=(0,1)$ become non-nodal arrivals to $Z=(-1,1)$ else only $(1-\eta)$ fraction of UAS lost become non-nodal arrivals. For any given $\lambda$, as more number of UAS are available to enter service when $\eta=0$, hence in $Z=(-1,1)$ we have more number of UAS in service, resulting in increased congestion, and more number of UAS in the queue when compared to $\eta=0.5$.
Figure 24: Theoretical and simulation comparison of congestion for $L=5,\ S=2L+1=11$ and varying arrival rate $\lambda\in[0.05,1]$, in a) zone $Z=(0,1)$ when the minimum number of UAS in the zone for congestion $M\in[2,4,6,8,10]$; and in b, c) zone $Z=(-1,1)$ when $M\in[2,3,4,5,6]$ and $\eta=0.5$ and $\eta=0$.
In Fig. 25, the expected traffic stream spread at level $1$, for given $M=2$, and varying arrival $\lambda$ can be determined from the average number of UAS in the service. For example, when $\lambda=0.8$ in Fig. 25a, the $U^{(1)}(1)\geq 1$ for $X\in\{-4,-3,-2,-1,0\}$, hence the expected traffic spread is in Stream$[X_{min}\leq X\leq X_{max}]$ segments, where $X_{min}=-4$ and $X_{max}=0$. As $\eta=0.1$, we have denser traffic flow on the left of the nominal segment, hence more spread towards $X<0$ streams. In Fig. 25b, $\eta=0.3$ and $U^{(1)}(1)\geq 1$ for $X\in\{-3,-2,-1,0,1\}$. Hence, the expected spread is in Stream$[-3\leq X\leq 1]$ segments. In Fig. 25c, as $\eta=(1-\eta)=0.5$, the $U^{(1)}(1)\geq 1$ for $X\in\{-2,-1,0,1,2\}$. The expected spread is in Stream$[-2\leq X\leq 2]$ segments; that is, we have equal spread distribution on the left and the right of the nominal segment.
@@
5.1 Presence of exogenous traffic stream
Following Assumptions 1-3, an exogenous traffic stream orthogonally intersecting the nominal segment at level $2$ is considered (refer Fig. 26a). The exogenous UAS arrival rate is $\lambda_{e}=0.2$. The expected number of exogenous UAS in the level $2$ zones is $S\lambda_{e}=2.2$. As $M$ must be greater than $2.2$ for UAS to cross the exogenous traffic stream, $M$ is chosen to be $4$. The exogenous UAS transitions with a nominal velocity of $1$ cell per timeslot. The exogenous UAS do not follow the proposed mitigation strategy and are hence unaffected by congestion in the zone. However, the exogenous UAS contributes to congestion. When $M$ or more than $M$ look-ahead positions (exogenous UAS included) are present in a zone, then the UAS laterally transitions to avoid such zone (as shown in snapshot 1 of Fig. 26a).
@@ 5.1 Presence of exogenous traffic stream
Unlike in the presence of a static obstacle, where it is certain that a UAS in $\beta\rightarrow\gamma$ path would always laterally transition every timeslot, here in the presence of an exogenous stream, the lateral transition of UAS is conditional (congestion-dependent). The $L$-slot look-ahead predicted positions of these UAS cannot be determined and are excluded in Definition $2$. This results in the UAS transitioning laterally and those transitioning upstream experiencing conflicts (as shown in snapshot 1-2 of Fig. 26a). A conflict among UAS belonging to the same stream is referred to as internal conflict. A conflict between a UAS and an exogenous UAS is referred to as exogenous conflict. Though conflicts occur, the congestion mitigation strategy guarantees no more than $M$ conflicts in any zone, that is UAS intrinsic safety is improved. Further, the mitigation strategy helps reduce the total number of conflicts in the stream (as shown in Fig. 26b). These advantages come at the expense of additional time delay in the UAS path and energy cost proportional to the number of diagonal transitions (refer Fig. 26b). Here, the total number of upstream and diagonal transitions executed to reach Level $5$ zone is $555\times S=$ 55). Note that the additional cost is insignificant when computed over the entire flight path. Depending on the inter-separation distance distribution between the exogenous UAS, the traffic spread may increase or decrease as shown in snapshots 3-4 of Fig. 26a. However, the queueing models can be used to estimate the expected traffic spread as validated in Fig. 26c.
@@
Figure 26: Congestion mitigation being employed by UAS in the presence of an exogenous traffic stream when $M=4$ and $\eta=0.3$. a) In snapshot 1, UAS 1 conflicts with UAS 2 belonging to the same stream at timeslot 47, and UAS 3 would conflict with an exogenous UAS 4 at timeslot 49. Rule 6 is employed by UAS to resolve the conflict. Snapshot 2 shows the UAS positions after the conflict is resolved. b) Comparison of the number of conflicts with ($M=4$) and without ($M=S$) the proposed congestion mitigation strategy and trade-off concerning time-delay and energy cost associated with the maneuver. c) Validation of expected traffic stream spread estimation in the presence of an exogenous stream ($\lambda_{e}=0.2$).
6 Conclusion